OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line?

OpenStudy (anonymous):

y= -2x +3

OpenStudy (anonymous):

these are the answers i can chose from y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3